OpenStudy (anonymous):

determine which ordered pair is a solution of y=-5+10 a(-15,5) b(2,0) c(-1,-5) d(3,25)

OpenStudy (michele_laino):

I think that your equation is: \(y=-5x+10\), am I right?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

in that case, if I replace \(x=2\) into my formula, I get: \[\huge y = \left( { - 5 \cdot 2} \right) + 10 = ...?\] please complete

OpenStudy (anonymous):

which solution

OpenStudy (michele_laino):

what is the value of \(y\) ?

OpenStudy (michele_laino):

please continue my computation above

OpenStudy (michele_laino):

hint: what is: \[\huge \left( { - 5 \cdot 2} \right) + 10 = ...?\]